QUESTION:
A town committee has a budget of $75 to spend on snacks for volunteers participating in a clean-up day. The committee chairperson decides to purchase granola bars and at least 50 bottles of water. Granola bars cost $.50 each, and bottles of water cost $.75 each. Write and graph a system of linear inequalities for the number of bottles of water and the number of granola bars that can be purchased.
I don't know where to start!!!

Answer by richwmiller(17219)

50g+75w<=7500
g>=50
w<=-2/3(g-150)
50*75=3750 50 bottles of water will cost 37 50 leaving 3750 for granola and water
50g+75w<=3750
You can get 75 granola bars or 50 more water bottles